### v1

# H3AE: High Compression, High Speed, and High Quality AutoEncoder for Video Diffusion Models

Abstract
Autoencoder (AE) is the key to the success of latent diffusion models for image and video generation, reducing the denoising resolution and improving efficiency. However, the power of AE has long been underexplored in terms of network design, compression ratio, and training strategy. In this work, we systematically examine the architecture design choices and optimize the computation distribution to obtain a series of efficient and high-compression video AEs that can decode in real time on mobile devices. We also unify the design of plain Autoencoder and image-conditioned I2V VAE, achieving multifunctionality in a single network. In addition, we find that the widely adopted discriminative losses, i.e., GAN, LPIPS, and DWT losses, provide no significant improvements when training AEs at scale. We propose a novel latent consistency loss that does not require complicated discriminator design or hyperparameter tuning, but provides stable improvements in reconstruction quality. Our AE achieves an ultra-high compression ratio and real-time decoding speed on mobile while outperforming prior art in terms of reconstruction metrics by a large margin. We finally validate our AE by training a DiT on its latent space and demonstrate fast, high-quality text-to-video generation capability.
**footnotetext: Equal contribution.
1 Introduction
Over the past year, diffusion models have revolutionized content creation by significantly improving generation quality and expanding text-to-image (T2I) generation [47, 45, 12, 4, 3, 2] to the more complex text-to-video (T2V) task [6, 71, 46, 40]. Recent advances from both industry [46, 59, 6] and the research community [65, 71, 55] now enable the creation of cinematic videos from text prompts [40, 59] or image inputs [5, 46]. These models also empower broader applications such as video editing [25, 36] and 3D generation [60, 33].
Figure 1: Compression ratio is in $log_{2}$ scale. Our H3AE not only achieves better compression ratio vs. PSNR tradeoff, but is also faster and parameter efficient. Refer Tab. 5 for more benchmarks.
The main powerhouse behind this success is Latent Diffusion Modeling (LDM) [47, 39]. Unlike pixel diffusion models [41, 50, 21, 10] which learns the diffusion mapping from noise to raw pixel space, LDMs learn the diffusion process in a compressed latent representation provided by a Variational Autoencoder (VAE) [30, 66, 1], which results in significantly better training and inference efficiency. Quite a few work have been done to improve the generation efficiency and mobile deployment of LDMs, which either compresses the denoising network [34, 23, 70, 62], or reduces the denoising steps [32, 64, 68]. However, the potential of VAEs has long been underexplored from the efficiency perspective. Intuitively, for a transformer denoiser, reducing the size of the latent space by a factor of $O(n)$ would result in theoretical gains of up to $O(n^{2})$ in computational cost during inference. This speed up is of utmost importance to video generation tasks considering the extra temporal dimension. Furthermore, as mentioned in [34, 62], VAE decoding emerges as a significant speed bottleneck in the generation process after successfully optimizing the denoiser by reducing its computation and steps, posing the demand for making VAE efficient while preserving reconstruction quality.
Previous work [9, 1, 18] has explored various aspects of VAE design in an ad-hoc manner, without a comprehensive study on architecture choices and training strategies. Key open questions include: (i) Compression: VAE compression significantly boosts the speed of diffusion models, especially for video generation. What is the highest achievable VAE compression ratio and how can we obtain better Pareto optimality for compression vs. quality? (ii) Efficiency: Given that VAE operates at pixel space, and thus is at risk of running slow or overflowing memory, how to design the network architecture to best balance performance and efficiency? (iii) Quality: How to effectively train video VAEs, i.e., what kind of augmentations and auxiliary losses can enhance their reconstruction quality? In this work, we provide a comprehensive study to address these questions.
Optimize for High Compression & High Speed VAE. We perform an ablation study on design choices (normalization, upsampling, latent channels, etc.) to decide the best VAE architecture. We find that 3D Causal Convolution in the entire network, especially in the high resolution stages, is neither efficient nor necessary for reconstruction quality. Instead, we use 2D convolutions to form disentangled spatial stages to balance quality and efficiency. We show that this is empirically better than using a non-parameterized patchify layer as in LTX [18]. Decoupling spatial stages allows us to slice and perform frame-wise encoding and decoding at high resolution stages. We further introduce 3D Causal Attention in video Autoencoder to achieve better performance with even less parameters. With the optimized design space, we perform mobile latency profiling and distribute the computation accordingly to form an efficient VAE backbone.
Training Algorithm to Improve Quality. Our proposed novel training algorithm consists of the following two components:
  • (a) Omni-Objective. Recently, Reducio [57] proposed a VAE design for image-to-video generation by leaking the first frame to the decoder to improve reconstruction quality with a higher compression ratio compared to T2V. To leverage this advantage and still support a single VAE for T2V and I2V generations, we propose a novel design that unifies decoding irrespective of the presence of image condition. We randomly drop the hierarchical features of the first frame as a condition to the decoder during training, such that our VAE can optionally act as an I2V VAE with improved quality. Surprisingly, this training strategy improves reconstruction even without image conditioning (plain T2V setting) due to the effect of training augmentation.
  • (b) Latent Consistency Loss. Existing works [65, 31, 18, 61] mainly use L1 reconstruction loss, KL regularization, LPIPS loss, and GAN loss to train the VAE. We show that these discriminative losses provide little performance improvements but significantly slow down training. We propose a new training strategy where we first train only with reconstruction loss and KL regularization for faster convergence. Then, we utilize a novel latent consistency loss to further finetune the model, by re-encoding the reconstructed video to obtain the fake posterior and optimize its KL divergence against the former posterior encoded from the ground truth video. This simple loss does not require complicated discriminator design but improves video reconstruction fidelity, both qualitatively and quantitatively.
Our contributions can be summarized as follows,
  • We systematically investigate the design of video VAEs, and propose an efficient architecture that achieves high compression ratio and high inference speed, while maintaining high reconstruction quality.
  • We propose to inject and randomly dropout the first frame condition to train a single VAE supporting both T2V and I2V tasks. The proposed method not only gives us a stronger I2V VAE, but also boosts the reconstruction quality in the pure T2V setting.
  • We revisit VAE training, proposing a novel, simple, and stable latent consistency loss to boost the reconstruction quality, outperforming all prior losses such as LPIPS, GAN and DWT.
  • We validate our VAE by training a 2B DiT in its highly compressed latent space, and demonstrate the first ever DiT-based video generation on mobile device.
Figure 2: Overview of H3AE architecture, omni-objective training, and Latent Consistency Loss. When computing $z^{\prime}$ in Eq. 2, the encoder weights remain frozen. For omni-objective training, we randomly pass the hierarchical features of the first frame from the encoder to the decoder, and use addition by default for feature fusion. As in the right, a block-shaped causal mask is applied to the 3D Transformer to enforce the causality of the attention mechanism, ensuring proper temporal dependencies in the generated representations.
2 Related Work
Latent Diffusion Models. Diffusion models [58, 47, 52, 20, 43, 29] have emerged as the dominant framework for generative modeling, surpassing traditional approaches like GANs [15, 28] and VAEs [30]. These models learn to iteratively denoise an initial noise sample into semantically meaningful and rich content through user defined input controls such as text-prompt and image. To learn large-scale diffusion models, early attempts such as text-to-image (T2I) [50, 10] and text-to-video(T2V) [41, 21] involved directly mapping the noise to the raw pixel space. While these works demonstrated the potential of diffusion models in generative modeling, running very deep neural networks in raw pixel space is computationally very expensive. This led to the development of Latent Diffusion Models (LDMs) [47, 5], which first learn a bi-directional mapping to compress the raw pixel space into a compressed and semantically rich latent space, followed by learning a diffusion model in this latent space. This helps in reducing the computational cost.
Many state-of-the-art T2I [45, 12, 4, 27, 14, 56, 38, 34, 23] and T2V [6, 71, 41, 46, 18, 65] models follow this latent diffusion modeling paradigm to synthesize visually stunning and semantically aligned content. For instance, SDXL [45] trains a large UNet architecture in the latent space to learn a foundational T2I model. SD3 [12] and Flux [4] improves upon this model by training a purely transformer model in a much more advanced latent space and changing the diffusion process to rectified flow for improved generation quality and faster sampling speed. Similarly, many large foundational T2V models utilize LDM framework for video generation. These include Stable Video Diffusion [5], Sora [6], Mochi [55], CogVideoX [22, 65], MovieGen [46] etc. In SnapGen-V [62], the compression factor of the latent models and reduced denoiser foot-print helps to enable their deployment on mobile devices.
Video Autoencoders. Autoencoders play a crucial role in determining both the quality and efficiency of LDMs, leading the research community to dedicate significant effort to improving their reconstruction performance and compression rates. Early foundational video LDMs [5, 17] were recycling image-based $8{\times}8$-compression AEs for the video domain. Such design was leaving a lot of temporal redundancy in the latent representations, and, to support longer video generation, later works (e.g., [65, 71, 72, 31]) pursued $4{\times}8{\times}8$ compression with the causal temporal structure [66]. Most recent foundational video LDMs explore AEs with higher compression rates of $8{\times}8{\times}$ 8 × 8 [46, 1, 54] or even $8{\times}16{\times}16$ [40]. While modern video AEs follow a typical convolutional design [65, 49, 69], several works explored wavelet transforms [16] to efficiently reduce the expensive representation dimensionality of high resolutions [37, 1, 35]. Other works develop transformer-based architectures for video auto-encoding [13, 9, 19, 67, 8].
[…]
3 Method
Spatial Upsample $|z|$ Norm Attention Params (M) FPS@512×512@$\times 512$ 512 PSNR SSIM rFVD
2D PixelShuffle 256 PN 196 38.1 29.48 0.8280 147.1
3D PixelShuffle 256 PN 189 29.53 0.8255 126.4
Patchify PixelShuffle 256 PN 194 43.5 28.46 0.7990 195.3
2D Interpolation 256 PN 169 40.1 24.78 0.7277 995.8
2D PixelShuffle 128 PN 195 38.3 28.34 0.7963 300.9
2D PixelShuffle 256 GN 197 37.9 29.43 0.8242 151.7
2D PixelShuffle 256 LN 197 27.5 29.49 0.8288 144.8
2D PixelShuffle 256 PN 316 41.7 28.59 0.8052 246.7
Table 1: AE Design Ablation. We ablate on various design choices for autoencoder architecture. The decoding latencies are benchmarked on iPhone 16 PM with 17 frames $512\times 512$ resolution output and the reconstruction quality evaluations are conducted on DAVIS dataset.
3.1 VAE Architecture
3.1.1 Micro Design
Inspired by most recent video VAEs [65, 1, 18], we start from a plain autoencoder backbone built with 3D Causal Convolutions, and explore the following design choices to construct a powerful yet efficient architecture.
Disentangled Spatial Stages.  Applying Conv3D at every stage of a video autoencoder can result in excessive memory consumption, especially in high-resolution stages, thereby hindering its efficient mobile deployment. To mitigate this issue, we disentangle the autoencoder structure by using 2D spatial convolutions in high-resolution stages, as illustrated in Fig. 2. This disentanglement reduces computation and peak memory for high-resolution stages and, in addition, allows chunked inference by slicing video frames. Notably, disentangling the high-resolution stage with Conv2D only has a negligible impact on the reconstruction quality compared to Conv3D, as shown in Tab. 1, while the non-parameterized patchify introduced by LTX [18] demonstrates degraded performance.
Upsampling. Besides traditional interpolation methods, PixelShuffle has become a popular trend in Autoencoders [9]. It requires additional learnable parameters to expand the feature channels, and performs channel-to-space/time operations as upsampling. We observe that PixelShuffle gives much better reconstruction quality than interpolation, with negligible overhead in parameters and latency, as in Tab. 1.
Latent Channels.  Prior continuous variational autoencoders typically adopt a small number of latent channels, such as $4ch$ in stable diffusion [48]. However, recent work have shown that using more latent channels not only improves reconstruction quality, but also offers better semantic completeness and aesthetic quality for diffusion models [12, 22, 23]. We explore larger number of latent channels for our VAE (i.e. 128-channel and 256-channel) in Tab. 1. We find that both 128-channel and 256-channel give reasonable reconstruction performance, and the 256-channel one achieves relatively higher PSNR. We report video diffusion results for both settings.
Normalization Layers.  Though static BatchNorm is foldable during the inference phase thus most efficient, we find that it fails to generalize to different resolutions for VAE reconstruction. We explore dynamic norms including PixelNorm [28] (PN), LayerNorm (LN) and GroupNorm (GN) in this work. Among them, LN is the default choice in transformer-based models, while GN is popular in CNN-based generative models [22, 65]. We find that PN, LN, and GN have similar reconstruction performance, as in Tab. 1. Note that PN is the simplest without learnable parameters, and in addition, LN and GN require reshape operations to process 5D video data, which are less efficient on mobile devices. As a result, we choose PN as our normalization method.
3.1.2 3D Causal Attention
With the operator selection optimized, we introduce 3D Causal Attention as an alternative foundation block in our VAE. 3D Causal Attention has a global spatial receptive field, while only attends to present and previous frames in temporal dimension. Tokens are flattened from $\mathbb{R}^{T\times H\times W\times C}$ to $\mathbb{R}^{(T\cdot H\cdot W)\times C}$ before being processed by the attention mechanism. To enforce causality, a block causal mask is applied to the attention score, where each block has the size of $L_{H}\times L_{W}$ indicating that the tokens are within a single frame, as illustrated in Fig. 2 (right). The mask ensures that video tokens from future frames remain inaccessible to tokens from earlier frames, which preserves temporal consistency and enables arbitrary-length video processing. In addition, QK-norm and RoPE [53] are also incorporated to improve training stability and convergence speed. Notably, RoPE also enhances the generalization capability to various resolutions, which is crucial for a foundational video Autoencoder. As in Tab. 1, using Causal 3D Attention reduces $38\%$ parameters, but achieves 0.9 higher PSNR.
3.1.3 Efficient and High Compression Autoencoder
Wrapping up the design, we divide both the encoder and the decoder into two stages. The high-resolution stage utilizes spatial Conv2D for sliced inference, the spatial-temporal stage uses 3D Causal convolution blocks as well as 3D Causal Attention to model temporal dependency and capture global receptive field, as in Fig. 2. Note that even though multiple downsampling (upsampling) sub-stages are involved in the spatial-temporal stage, we only distribute 3D Causal Attention in the bottleneck position with the highest compression ratio. This straightforward design aligns the computation and memory complexity of the 3D Attentions in VAE with those in the DiT denoiser, ensuring that the entire diffusion pipeline is runnable on resource-constrained devices. For instance, our $8\times 32\times 32$ H3AE has only $1/32$ tokens in the bottleneck and latent stage compared to the CogVideoX $4\times 8\times 8$ VAE, thus the 3D Attentions in our H3AE and our DiT enjoy more than 1000×1000$\times$1000 × FLOPs reduction compared to CogVideoX DiT, significantly improving computational efficiency for video generation.
Besides architecture design, it is also crucial to properly distribute network parameters and computations to achieve a better Pareto curve of performance and efficiency. We investigate the depth and width scaling property of our VAE by directly profiling on the mobile device (iPhone 16 PM). We intuitively set the maximum network width based on the memory bound, and set a real-time target for mobile inference. i.e., decoding a $17\times 512\times 512$ video clip within 0.5s, which is equivalently more than 30 FPS. We scale the network by shrinking the width but increasing the depth to maintain this real-time target, and explore reconstruction performance, as in Tab. 2. We find that in this scope, wide but shallow VAE achieves the best reconstruction quality. All of our H3AEs are then constructed based on depth-width ratios of the 1×1$\times$1 × variant in Tab. 2.
Width Params (M) Latency (s) PSNR
0.5×0.5$\times$0.5 × 233.0 0.45 27.02
0.75×0.75$\times$0.75 × 206.1 0.46 29.11
1×1$\times$1 × 196.5 0.45 29.48
Table 2: Scaling of network backbone. The 1×1$\times$1 × backbone is constructed under real-time constraint while maximizing the width to fit mobile memory.
[…]
$8\times 32\times 32$ 29.48 0.8280 0.2627 0.2611 147.1
omni-plain 29.95 0.8367 0.2506 0.2470 129.6
omni-I2V 30.08 0.8384 0.2456 0.2427 128.4
$8\times 64\times 64$ 25.21 0.7284 0.3937 0.4048 737.9
omni-plain 25.97 0.7387 0.3963 0.4041 709.7
omni-I2V 26.38 0.7503 0.3748 0.3879 684.1
Table 3: Omni-AE for the T2V and I2V tasks. The baseline is trained without image condition while omni-training utilized the image condition with probability ($p=0.5$). The result shows that adding the omni-training objective helps both plain reconstruction and image-conditioned setting for I2V.
3.2 Omni-AE for both T2V and I2V
Image-to-video (I2V) generation is a prominent application in video generation [65, 71, 57], aiming to synthesize realistic and temporally coherent video sequences conditioned with a user-specified input image. This task is crucial for various applications, including video prediction, animation, content creation, and scene understanding.
In the I2V task, an image (typically the first frame) is provided as the generation condition. Tian et al. [57] propose to construct an image-conditioned VAE to utilize this additional information, achieving better decoding quality in high compression settings. Though utilizing the image condition benefits quality, the dedicated nature of I2V-VAE prohibits its wide deployment. Considering the cost of adapting the denoiser to a new VAE when swapping from T2V to I2V, most video diffusion works [5, 22, 46] still use plain VAEs for simplicity.
In this work, we propose a simple but effective multifunctional VAE that works for both plain T2V and conditioned I2V setting. Specifically, we propose an omni-objective training, where we take the hierarchical features of the first frame from the encoder and feed them to the decoder as condition with probability $p$. The decoder simulates I2V scenario when receiving the condition features, while performing plain reconstruction when not. As shown in Tab. 3, we find that after the omni-objective training, our VAE can perform both plain reconstruction and image-conditioned reconstruction effectively with a single set of weights. Interestingly, because of the auxiliary information, not only does the I2V setting get enhanced quality, but we also find that this training method improves plain reconstruction. This finding shows that it is a free lunch to train Omni-AEs serving both tasks: plain reconstruction and I2V setting. For the main experimental results in the following section, unless otherwise stated, we train H3AEs with this omni-objective and inference with the plain reconstruction setting to fairly compare with baselines.
[…]
Baseline 29.48 0.8280 0.2627 0.2611 147.1
LPIPS 29.07 0.8205 0.224 0.2275 148.9
GAN 29.35 0.8241 0.2631 0.2623 161.0
DWT 29.36 0.8247 0.2651 0.2642 159.3
Ours 29.58 0.8299 0.2541 0.2511 142.2
Table 4: Training loss comparison. Baseline is $L_{\text{recon}}+\lambda L_{\text{KL}}$. Ours represent the baseline followed by fine-tuning with the latent consistency loss. It shows that our proposed loss function yields better reconstruction quality (measured by rFVD) and achieves considerably better metrics than other auxiliary losses.
Model $f$Params. (M) FPS@512×512@$\times 512\uparrow$ 512 ↑ DAVIS OpenVid-HD
[…]
CogVideoX-VAE $4\times 8\times 8$ 205.60 13.6 32.37 0.8954 26.30 37.88 0.9713 0.81
Hunyuan-VAE $4\times 8\times 8$ 235.06 15.5 32.91 0.8951 21.17 39.07 0.9738 0.51
CV-VAE $4\times 8\times 8$ 173.99 24.9 27.13 0.7722 220.61 31.78 0.9095 6.04
OpenSora-1.2-VAE $4\times 8\times 8$ 375.12 21.7 14.2 28.76 0.8091 193.42 32.98 0.9192 6.93
Cosmos-CV $4\times 8\times 8$ 70.74 141.7 40.5 31.20 0.8654 34.58 35.16 0.9492 1.95
Cosmos-CV $8\times 8\times 8$ 107.26 130.8 42.8 29.34 0.8233 89.27 33.84 0.9371 2.51
Cosmos-CV $8\times 16\times 16$ 107.26 161.9 62.0 26.68 0.7558 319.21 30.74 0.8916 15.56
LTX-VAE $8\times 32\times 32$ 399.77 188.0 17.9 28.34 0.7923 165.43 34.62 0.9371 3.82
H3AE $4\times 16\times 16$ 67.16 269.8 78.3 32.96 0.8987 20.36 37.43 0.9666 0.99
H3AE $8\times 32\times 32$ 196.51 195.4 38.1 30.23 0.8412 122.82 35.23 0.9523 3.55
H3AE $8\times 64\times 64$ 643.76 182.8 39.0 26.03 0.7404 688.4 30.23 0.8788 39.3
Table 5: SOTA AE Comparison. Comparing with state-of-the-art autoencoders on latency, reconstruction quality on DAVIS and OpenVid datasets. For a fair comparison, we report our results without image condition (plain VAE). The I2V performance of our VAE is shown in Tab. 3 and Fig. 5. Note that our VAE is fully causal.
3.3 Training Losses
Current Variational Autoencoder (VAE) training commonly employs $\mathcal{L}_{1}$-norm as reconstruction loss ($\mathcal{L}_{\text{recon}}$), KL as regularization ($\mathcal{L}_{\text{KL}}$) on latents, and multiple auxiliary losses ($\mathcal{L}_{\text{aux}}$) to improve reconstruction performance. This results in the following total loss:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{recon}}+\lambda\mathcal{L}_{\text{KL}}+\mathcal{L}_{\text{aux}}$ (1)
Existing works have investigated various auxiliary losses such as Perceptual loss [26], Discrete Wavelet Transform (DWT) loss [37, 18], GAN loss [18, 31, 65], etc. Perceptual loss measures the reconstruction error in feature space of a pre-trained neural network. Similarly, DWT loss computes the feature difference in a wavelet frequency space, while GAN loss introduces a discriminator to learn the separation between real and fake samples. In our empirical study on large-scale video training, we find that these auxiliary losses hardly provide improvements over the baseline, as shown in Tab. 4. Plus, most of them substantially slow down the training because of the large computation graph, and brings visual artifacts as also noticed by [18].
3.3.1 Latent Consistency Loss
To overcome the drawbacks of these discriminative losses, we design a new auxiliary training loss to enhance reconstruction quality by utilizing the unique paradigm of VAEs. Specifically, we reuse the VAE encoder as the discriminator, and encode the reconstructed video to obtain a fake posterior ($z^{{}^{\prime}}$). Note that the VAE encoder weights are not updated at this step. We compare the KL Divergence between the fake posterior ($z^{{}^{\prime}}$) and the posterior ($z$) encoded from the ground truth video, which has already been computed in the training forward pass. Thus, the latent consistency loss (LC loss) is:
$\begin{split}\mathcal{L_{\text{LC}}}=&D_{KL}(z,z^{\prime})\\ =&\frac{1}{2}\sum\left(\frac{(\mu_{z}-\mu_{z^{\prime}})^{2}}{\sigma_{z^{\prime}}^{2}}+\frac{\sigma_{z}^{2}}{\sigma_{z^{\prime}}^{2}}-1-\log\frac{\sigma_{z}^{2}}{\sigma_{z^{\prime}}^{2}}\right)\end{split}$ (2)
This design holds several advantages. First, we eliminate the need to design complicated discriminators as in GAN loss, and the training is more stable without the risk of mode collapse. Second, the discriminative loss provides stronger guidance compared to LPIPS and DWT, as the encoder is inherited on the fly from the VAE itself, which seamlessly preserves the knowledge and evolves itself. As in Tab. 4, we show that our proposed LC loss consistently improves reconstruction.
Figure 3: AE Qualitative Results. Reconstructions from our H3AE ($8\times 32\times 32$) and other high compression autoencoders: Cosmos-Tokenizer [1] ($8\times 16\times 16$), LTX-VAE [18] ($8\times 32\times 32$). We show zoomed-in results to highlight the differences in fidelity and quality. Our method features greater high-frequency detail. Here ‘GT‘ refers to the ground truth video.
4 Experiments
[…]
GPU iPhone AQ IQ MS Quality Semantic Total
Cosmos-Tokenizer 16 $8\times 16\times 16$ 0.40 0.5606 0.5097 0.9875 0.8101 0.6860 0.7853
LTX-VAE 128 $8\times 32\times 32$ 0.09 1.00 0.5981 0.6028 0.9896 0.8230 0.7079 0.8000
H3AE 128 $8\times 32\times 32$ 0.09 1.00 0.6017 0.6073 0.9875 0.8342 0.7147 0.8103
H3AE 256 $8\times 32\times 32$ 0.09 1.00 0.6170 0.6314 0.9885 0.8338 0.7226 0.8110
H3AE 256 $8\times 64\times 64$ 0.05 0.22 0.5441 0.5317 0.9901 0.8003 0.6153 0.7633
Table 6: Quantitative Metrics on VBench. Text-to-video generation benchmark on Vbench [24] with a 2B DiT denoiser trained using different Autoencoders. Overall Vbench score and selected score are reported. We specifically exhibit aesthetic quality (AQ), imaging quality (IQ), and motion smoothness (MS) to evaluate the performance of denoisers. Latency of DiT is benchmarked by generating 65-frame $512\times 512$ video clips on an NVIDIA A100 80GB GPU and 17-frame $512\times 512$ clips on iPhone 16 Pro Max. Notably the DiT using Cosmos-Tokenizer results in OOM error on iPhone due to memory inefficiency.
Figure 4: T2V Qualitative Results. Examples of videos generated by a 2B DiT denoiser, trained on the latent space of our $8\times 32\times 32$ H3AE.
Implementation details.  The model is trained using our internal video dataset. We resize and randomly crop a $256\times 256$ patch from each video clip for training. The model is initially trained using 17-frame video clips for 70K iterations for faster convergence, and then finetuned on video clips with various lengths (1-65) for 10K iterations for better generalization. Latent consistency loss is only applied in the second stage. The training is conducted on 32 NVIDIA A100 80G GPUs, with a batch size of 16 per GPU. We use AdamW optimizer with $41e-$ 4 learning rate and $\beta=[0.9,0.999]$ for the first stage, and reduce the learning rate to $51e-$ 5 in the second stage.
Evaluation.  We benchmark reconstruction quality using high resolution video dataset, including 60 video clips from DAVIS-2017 [44] testset and 4000 high resolution video clips randomly sampled from OpenVid [42] (OpenVid data is excluded in our training to ensure zero-shot evaluation). The first 33 frames from each clip are utilized for evaluation with a spatial resolution at $512\times 512$. We evaluate the PSNR, SSIM, and reconstruction-FVD (rFVD) to benchmark the performance of autoencoders.
4.1 Comparison with SOTA autoencoders
We compare our model with SOTA video tokenizers, i.e. CogVideoX-VAE, CV-VAE [69], Cosmos-Tokenizer [1], LTX-VAE [18], etc. Among them, Cosmos-Tokenizer and LTX-VAE are focused on high compression ratios. Three variants of H3AE models are built up with $4\times 16\times 16$, $8\times 32\times 32$, and $8\times 64\times 64$ compression ratios respectively to demonstrate the robustness of our design choices and analysis. We obtain the pre-trained SOTA models and evaluate them under the same evaluation setting mentioned above for reconstruction quality. We report inference speed on Nvidia A100 GPU and iPhone 16 Pro.
As shown in Tab. 5 and Fig. 1, our models achieve better reconstruction under the same or higher compression ratio. Specifically, with 4×4$\times$4 × higher compression ratio, our $4\times 16\times 16$ VAE outperforms Cosmos-Tokenizer ($4\times 8\times 8$) across both evaluation sets, achiving notable improvements of $+1.76$ PSNR, $+0.0333$ SSIM, $-14.22$ rFVD on DAVIS, as well as and $+2.27$ PSNR, $+0.0174$ SSIM, $-0.96$ rFVD on OpenVid-HD. Furthermore, our $8\times 32\times 32$ VAE significantly outperforms LTX-VAE by $+1.89$ PSNR, $+0.048$ SSIM, $-42.61$ rFVD on DAVIS, along with $+0.61$ PSNR, $+0.0152$ SSIM, $-0.27$ rFVD on OpenVid-HD. In addition, our model demonstrates superior overall efficiency. Our $8\times 32\times 32$ VAE requires only half the parameters compared to LTX-VAE, and achieves 2.5×2.5$\times$2.5 × speedup faster on iPhone, making it significantly more lightweight and faster for real-time applications.
Visual Quality. We exhibit a visual comparison of reconstructed samples between Cosmos-Tokenizer ($8\times 16\times 16$), LTX-VAE, and our H3AE ($8\times 32\times 32$) in Fig. 3. The results demonstrate that our model delivers better contextual fidelity and high-frequency quality than Cosmos-Tokenizer. Noteworthy, the visual comparison also demonstrates the effectiveness of our proposed training method, while Cosmos-Tokenizer and LTX-VAE either introduces overly smooth reconstruction or unpleasant artifacts.
4.2 Video Generation
We show text-to-video generation results in Tab. 6 and Fig. 4 to demonstrate that our VAE creates a good and highly-compressed latent space for video diffusion models. We construct a DiT with 2B parameters aligned with popular video diffusion model settings. Specifically, we use 3D full attention with RoPE and QK-Norm in transformer blocks. All models are first trained on image dataset to learn spatial knowledge for $50K$ iterations, then these model are trained with image-video joint training strategy for $100K$ iterations. The VDM training is done on 128 A100 GPUs for 4-10 days depending on the VAE compression ratio. As shown in Tab. 6, we find that using our H3AE ($8\times 32\times 32$) outperforms Cosmos-Tokenizer ($8\times 16\times 16$) and LTX-VAE in Vbench score. Our high-compression ratio autoencoder also delivers fast and memory efficient inference on both GPU and edge device. We additionally report the result of using an extremely high-compression ratio ($8\times 64\times 64$) H3AE, which provides more than a 4×4$\times$4 × speed-up on the iPhone 16 PM. This demonstrates the great potential of our approach for efficient video generation on edge devices.
5 Conclusion
In this work, we systematically examine autoencoder architecture design and optimize the computation distribution to obtain a series of efficient AEs that can decode latents in real-time on mobile-device. With the ultra high spatial-temporal compression ratio, we successfully reduce the latent tokens and achieve faster generation speed for the DiT-based video diffusion model. We unify plain reconstruction and image-conditioned I2V reconstruction, demonstrate improved results for both settings with a single set of weight, simplifying applications and saving the potential adaptation cost. Our empirical study also shows that popular discriminative losses, i.e., GAN, LPIPS, and DWT losses, provide no significant improvement when training AEs at scale. We propose a novel latent consistency loss that requires no complicated discriminator design or hyperparameter tuning but provides stable improvements in reconstruction quality.
Appendix
Figure 5: Quality comparison of reconstruction results of H3AE between plain-T2V VAE and I2V VAE settings. The results shows that I2V VAE delivers better high-frequency details.
AE $f$ Params DAVIS OpenVid-HD
PSNR SSIM LPIPS PSNR SSIM LPIPS
DC-AE $32\times 32$ 308.4 28.28 0.7886 0.0732 30.31 0.8709 0.0474
H3AE $32\times 32$ 196.5 36.27 0.9475 0.0724 38.11 0.9694 0.0378
DC-AE $64\times 64$ 645.5 28.14 0.7865 0.0746 30.16 0.8689 0.0479
H3AE $64\times 64$ 643.8 29.50 0.8169 0.2640 31.04 0.8822 0.1695
Table 7: Despite H3AE is designed as video autoencoder, it can also be used as an image autoencoder. The quality comparison of image reconstruction between current SOTA high-compression autoencoder i.e. DC-AE[9] and our H3AE on DAVIS and OpenVid-HD datasets. The results demonstrate H3AE outperforms DC-AE under same compression ratios.
Figure 6: More examples of videos generated by a 2B DiT denoiser, trained on the latent space of our H3AE.